# SGD with Clipping is Secretly Estimating the Median Gradient

### v1

@@ 2.2 Mean Estimation and (Sample) Median
The choice of $\mathcal{D}$ determines properties of the associated estimator: if the function $\mathcal{D}$ heavily penalizes large values, like the quadratic $\mathcal{D}=\tfrac{1}{2}\|\cdot\|_{2}^{2}$, then the solution to (4) will be sensitive to outliers. In contrast $\mathcal{D}=\|\cdot\|_{2}$ and $\mathcal{D}=\|\cdot\|_{1}$ grow only linearly, and thus (4) will be less sensitive to outliers. We return to problem (4) later in Section 4.
For SGD, its update direction is a sampled gradient $\bm{g}_{t}$ which aims to approximate the full gradient $\nabla\ell(\bm{w}_{t})$. This section motivates that in the presence of large outliers it might be necessary to use more robust gradient estimators as update direction. A natural alternative is to use the sample median together with SGD (Alistarh et al., 2018; Acharya et al., 2022), as we do next.
3 Warm-up: Sample Median Gradient
@@
In this section, we showcase that SGD with an (approximate) sample median is robust to heavy-tailed noise, whereas SGD using the sample mean can fail. Consider the following setup: at every iteration $t\geq 0$, we have access to $n$ noisy oracle gradients $\big{(}\bm{g}_{t}^{(1)},\ldots,\bm{g}_{t}^{(n)}\big{)}$, and we apply
@@
where $\bm{m}_{t}=\texttt{median}_{\ell_{p},i\in[n]}(\bm{g}_{t}^{(i)})$, and $\bm{e}_{t}\in\mathbb{R}^{d}$ represents an error we may commit in computing the sample median.
@@
There exists $\sigma_{1},\sigma_{2},\delta_{1},\delta_{2}\geq 0$ for every $\bm{w}\in\mathbb{R}^{d}$, such that the sample median $\bm{m}=\texttt{median}_{\ell_{p},i\in[n]}(\bm{g}^{(i)})$ verifies
@@
Let $\ell:\mathbb{R}^{d}\to\mathbb{R}$ be $L$-smooth and let Assumption  3.1 hold with $\delta_{2}\leq\frac{1}{8}$. Take $\eta\leq\frac{1}{8L(1+2\sigma_{2}+2\delta_{2})}$ and consider $\bm{w}_{t}$ a sequence generated by (5) where $\bm{e}_{t}$ verifies Assumption 3.2. Then, for every $T\geq 1$ it holds
@@
Let $\ell:\mathbb{R}^{d}\to\mathbb{R}$ be convex and $G$-Lipschitz, with ${\rm{argmin}}~{}\ell\neq\emptyset$. Let Assumptions 3.1 and Assumption 3.2 hold. Without loss of generality assume that $\delta_{2}=\sigma_{2}=0$. Then, for $\eta=\frac{1}{(\delta+\varepsilon)T+\sqrt{T}}$ and every $T\geq 1,$ the iterates $\bm{w}_{t}$ from (5) satisfy
@@
The proofs are provided in Appendix A. Let us now discuss our main Assumption 3.1. Assume for simplicity that $\delta_{2}=\sigma_{2}=0$, that $\bm{e}_{t}=0$, and consider that $\bm{g}_{t}^{(i)}=\nabla\ell(\bm{w}_{t})+\bm{\xi}_{t}^{(i)}$, where $\bm{\xi}_{t}^{(i)}$ is a noise term. Then, Assumption 3.1 becomes [^1] See Lemma A.2 in the appendix for a proof.
@@
In other words, we need the second moment of the sample median of $\bm{\xi}_{t}^{(i)}$ to be uniformly bounded. We want to stress that our choice of using the median is crucial when the $\bm{\xi}_{t}^{(i)}$ come from a heavy-tailed distribution, such as a (univariate) standard $\alpha$-stable distribution, with $\alpha\in(1,2)$. Indeed, in this case we know that the variance of the sample median is finite for sufficiently large sample size $n$ (see (Bickel, 1967, Theorem 2.2) and (Nolan, 2020, Theorem 1.2)) [^2] We know more: The sample median of a parent distribution with median $\xi$ is asymptotically normal with location $\xi$ (Chu & Hotelling, 1955). For Cauchy distributions ($\alpha=1$), the sample median has a finite variance for $n\geq 5$ (Chu & Hotelling, 1955, Remark 1), while first and second moment of the sample mean are undefined.. On the contrary, if we were taking $\bm{m}_{t}$ as a more classical sample mean we would be in trouble, because the variance of the sample mean would be infinite [^3] This is due to the fact that, by definition of $\alpha$-stability, the sample mean of i.i.d sample follows the same $\alpha$-stable distribution (up to affine translation). everywhere. See Fig. 1 for an illustration of this phenomenon.
This simple discussion illustrates that Assumption 3.1 can be satisfied under mild conditions on the noise. We exploit this in the next corollary, which shows that the $\ell_{1}$-sample median can even accommodate state-dependent noise. Let us define first our model for such noise.
@@
As opposed to the following assumption that has been commonly made in the recent literature (see e.g., (Zhang et al., 2020b; Puchkin et al., 2023; Sadiev et al., 2023)) $\mathbb{E}\left[\|\bm{g}-\nabla\ell(\bm{w})\|^{p}\right]\leq\sigma^{2}$ for some $p<2$ where $\sigma^{2}>0$ cannot depend on $\bm{w}$, Assumption 3.5 can accommodate much stronger noise. To illustrate this, consider $d=1$ and $\ell(\bm{w})=\bm{w}^{2}/2$. Assumption 3.5 allows the noise model $\bm{\Sigma}(\bm{w})\bm{\zeta}$ with $\bm{\Sigma}(\bm{w})=\sqrt{c_{1}+c_{2}\bm{w}^{2}}$, which results in strong noise, since $\bm{\zeta}$ has infinite variance if $\alpha<2$, and this is amplified by $\bm{\Sigma}(\bm{w})$ when $\bm{w}$ is large. Such strong noise makes vanilla SGD diverge, as we will illustrate in our experiments.
@@
Let $n\geq 3$ be odd and $p=1$. Consider the iterates from (5) with $\bm{e}_{t}=0$, and suppose that $\bm{g}_{t}^{(i)}=\nabla\ell(\bm{w}_{t})+\bm{\Sigma}(\bm{w}_{t})\bm{\zeta}_{t}^{(i)}$ verifies Assumption 3.5.
@@
    If $\ell$ is smooth and $\eta=1/\sqrt{T}$ then we have that $\min_{0\leq t\leq T-1}\mathbb{E}\left[\|\nabla\ell(\bm{w}_{t})\|^{2}\right]=\mathcal{O}(1/\sqrt{T})$.
@@
    If $\ell$ is convex and Lipschitz and $\eta=1/\sqrt{T}$ then we have that $\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]=\mathcal{O}(1/\sqrt{T})$, where $\bar{\bm{w}}_{T}$ is defined as in Proposition 3.4.
@@
This result illustrates how powerful the sample median can be when used in gradient-based optimization. However, the method (5) has clear drawbacks: (i) it requires to draw multiple samples per iteration (multiple calls of the oracle), and (ii) we need to compute the sample median up to some tolerance. For instance, the geometric median can not be computed in closed form and there exists a vast body of literature on how to approximate it (Cohen et al., 2016).
We tackle (ii) in the next section by developing an efficient and inexpensive iterative method for approximately solving the general problem (4) which includes the sample median as a special case. We tackle issue (i) later in Section 5.

### v2

@@ 2.2 Mean Estimation and (Sample) Median
The choice of $\mathcal{D}$ determines properties of the associated estimator: if the function $\mathcal{D}$ heavily penalizes large values, like the quadratic $\mathcal{D}=\tfrac{1}{2}\|\cdot\|_{2}^{2}$, then the solution to (5) will be sensitive to outliers. In contrast $\mathcal{D}=\|\cdot\|_{2}$ and $\mathcal{D}=\|\cdot\|_{1}$ grow only linearly, and thus (5) will be less sensitive to outliers. We return to problem (5) later in Section˜3.2.
3 Robust Gradient Estimation with Stochastic Proximal Point
3.1 Warmup: Momentum as Online Mean Gradient Estimation
When dealing with noisy gradient oracles, a standard technique is to apply (heavy-ball) momentum (Polyak, 1964). For momentum coefficient $\beta\in[0,1)$, it is given by
$\begin{split}\bm{m}_{t+1}=\beta\bm{m}_{t}+(1-\beta)\bm{g}_{t},\\ \bm{w}_{t+1}=\bm{w}_{t}-\eta_{t}\bm{m}_{t+1}.\end{split}$ (SGD-M)
Momentum has been empirically shown to improve the performance of SGD in machine learning (Sutskever et al., 2013), and can be seen as a variance reduction technique (Gower et al., 2020).
It is easy to see that momentum performs an exponentially weighted average over past gradients. In this section, we will derive a new perspective on SGD-M, namely being an online estimator of the mean gradient at a given point. To see this, we recall that for a random variable $\bm{z}\in\mathbb{R}^{d}$, its mean (if it exists) is the solution to (5) with $\mathcal{D}=\frac{1}{2}\|\cdot\|^{2}$. The random variable of interest in this context are the noisy gradients $\bm{g}_{t}$. We could in principle solve problem (5) with iterative stochastic methods. For this purpose, we consider the stochastic proximal point (SPP) method (Asi & Duchi, 2019; Davis & Drusvyatskiy, 2019). We give a detailed introduction to SPP in the next section.
For the purpose of this section, it is sufficient to know that the iterates $(\bm{m}_{t})$ of SPP with step-size $\tau>0$ applied to (5) are given by (details in Appendix B.2)
$\bm{m}_{t+1}\;=\;\underset{\bm{m}}{\operatorname{argmin}}\;\mathcal{D}(\bm{m}-\bm{g}_{t})+\frac{1}{2\tau}\|\bm{m}-\bm{m}_{t}\|^{2}=\bm{g}_{t}+\mathrm{prox}_{\tau\mathcal{D}}(\bm{m}_{t}-\bm{g}_{t}),$ (6)
$\mathrm{prox}_{\mathcal{D}}(\bm{x}):=\underset{\bm{y}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\mathcal{D}(\bm{y})+\frac{1}{2}\|\bm{y}-\bm{x}\|^{2}.$
For $\mathcal{D}=\frac{1}{2}\|\cdot\|^{2}$, it is easy to compute $\mathrm{prox}_{\tau\mathcal{D}}(\bm{x})=\frac{1}{1+\tau}\bm{x}$. Thus, one iteration of SPP (6) is equal to
$\bm{m}_{t+1}\;=\;\bm{g}_{t}+\mathrm{prox}_{\tau\mathcal{D}}(\bm{m}_{t}-\bm{g}_{t})=\bm{g}_{t}+\tfrac{1}{1+\tau}(\bm{m}_{t}-\bm{g}_{t})=(1-\tfrac{\tau}{1+\tau})\bm{m}_{t}+\tfrac{\tau}{1+\tau}\bm{g}_{t}.$ (7)
This is exactly the momentum step in (SGD-M) with $\beta=1-\tfrac{\tau}{1+\tau}$. However, in SGD-M the parameters $\bm{w}_{t}$ are updated each step as well, and thereby the distribution of the gradients $\bm{g}_{t}$ changes each step. In conclusion, the momentum operation can be understood as an online estimation of the mean gradient.
We end this section by justifying our choice to specifically consider SPP for solving (5), instead of SGD or other stochastic methods. If we were to apply SGD with step size to problem (5), we would obtain
$\bm{m}_{t+1}\;=(1-\tau)\bm{m}_{t}+\tau\bm{g}_{t}.$
This also resembles momentum, but note that the momentum coefficient $1-\tau$ could be negative if $\tau>1$. For SPP on the other hand, the momentum coefficient is correctly parametrized since $1-\tfrac{\tau}{1+\tau}$ is contained in $[0,1)$ for every $\tau>0$.
In conclusion, we can understand heavy-ball momentum (SGD-M) as applying SPP to the mean estimation problem in an online fashion. This opens the question what is the SPP update for estimators that are more robust to heavy-tailed noise, such as the median (or more generally for other choices for $\mathcal{D}$ that are of interest). We answer this in the next section.
3.2 Clipping as Online Median Gradient Estimation
We now focus on the following gradient estimation problem: let $\mathcal{G}$ denote the (fixed) distribution over noisy gradients $\bm{g}$. For $\mathcal{D}:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ being a closed, proper, convex function, we consider the problem of approximating
We consider stochastic iterative methods for solving (8), with a special focus on the stochastic proximal point method (SPP). We recall the SPP update being
$\begin{split}\bm{m}_{t+1}\;&=\;\underset{\bm{m}}{\operatorname{argmin}}\;\mathcal{D}(\bm{m}-\bm{g}_{t})+\frac{1}{2\tau}\|\bm{m}-\bm{m}_{t}\|^{2}\\ &=\bm{g}_{t}+\mathrm{prox}_{\tau\mathcal{D}}(\bm{m}_{t}-\bm{g}_{t}).\end{split}$ (9)
The SPP update can be seen as an implicit version of SGD, which is hard to compute in general. In cases where SPP has a closed form update, it is often preferred to SGD since is easier to tune (Asi & Duchi, 2019; Milzarek et al., 2024). Fortunately, as we show next, the SPP method enjoys closed-form updates for all choices of $\mathcal{D}$ which are of interest in our context, in particular for $\mathcal{D}\in\{\|\cdot\|_{1},\|\cdot\|_{2},\tfrac{1}{2}\|\cdot\|^{2}\}$. All the details are deferred to Section˜B.3.
Vectorwise clipping.
If we choose $\mathcal{D}=\|\cdot\|_{2}$, the SPP update applies clipping to the difference between sampled gradient $\bm{g}_{t}$ and previous estimate $\bm{m}_{t}$.
Corollary 3.1.
For $\mathcal{D}=\|\cdot\|_{2}$ update (9) is given by
$\bm{m}_{t+1}$ $=\bm{m}_{t}+\mathrm{clip}_{\tau,2}(\bm{g}_{t}-\bm{m}_{t}),$ (10)
We can rewrite update (10) as $\bm{m}_{t+1}=\beta_{t}\bm{m}_{t}+(1-\beta_{t})\bm{g}_{t}$ with $\beta_{t}:=1-\tfrac{\tau}{\max\{\tau,\|\bm{g}_{t}-\bm{m}_{t}\|_{2}\}}$. In other words, online estimation of the geometric median ($\mathcal{D}=\|\cdot\|_{2}$) with SPP is equal to a cautious version of momentum, as samples which are too far from $\bm{m}_{t}$ will be down weighted. This allows the iterates to be more robust to outliers.
Componentwise clipping.
If we choose $\mathcal{D}=\|\cdot\|_{1}$, SPP performs componentwise clipping instead. Here as well, clipping protects against large individual entries by shrinking them independently.
Corollary 3.2.
For $\mathcal{D}=\|\cdot\|_{1}$ update (9) is given by
$\bm{m}_{t+1}=\bm{m}_{t}+\mathrm{clip}_{\tau,1}(\bm{g}_{t}-\bm{m}_{t}),$ (11)
Huber function.
For update (10), it is possible that $\bm{m}_{t+1}=\bm{g}_{t}$, thus no momentum. To allow for a soft transition, we can choose $\mathcal{D}$ to be the Huber function. For $\mu>0$, it is given by
$H_{\mu}:\mathbb{R}^{d}\to\mathbb{R},\quad H_{\mu}(\bm{z})=\begin{cases}\frac{1}{2}\|\bm{z}\|^{2}&\quad\|\bm{z}\|\leq\mu,\\ \mu\|\bm{z}\|-\frac{\mu^{2}}{2}&\quad\text{else}.\end{cases}$
The Huber function operates like the squared loss for arguments with small norm, and like the $\ell_{2}$-norm for arguments with large norms (potentially outliers).
Corollary 3.3.
For $\mathcal{D}=H_{\mu}$, update (9) is given by
$\bm{m}_{t+1}$ $=\bm{g}_{t}+\beta_{t}(\bm{m}_{t}-\bm{g}_{t})=\beta_{t}\bm{m}_{t}+(1-\beta_{t})\bm{g}_{t},$ (12)
where $\beta_{t}:=1-\tfrac{\mu\tau}{\max\{\|\bm{m}_{t}-\bm{g}_{t}\|,~\mu(1+\tau)\}}$.
In update (12) the coefficient $\beta_{t}$ is always positive. Thus, $\mathcal{D}=H_{\mu}$ is a tradeoff between momentum with fixed coefficient ($\mathcal{D}=\frac{1}{2}\|\cdot\|^{2}$) and clipping ($\mathcal{D}=\|\cdot\|_{2}$). However, it comes at the cost of choosing/tuning an additional hyperparameter $\mu$. In Section˜B.4, we additionally compare the updates when using stochastic subgradient descent instead of SPP for solving (8), showing that sign-SGD is related to $\ell_{1}$-median estimation.
4 Gradient Estimator in the Wild
[…]
With the results from Sections˜3.1 and 3.2, we immediately obtain the following special cases:
(ii)  if $\mathcal{D}=\|\cdot\|_{p}$, $p\in\{1,2\}$ : $\begin{cases}\bm{m}_{t+1}=\bm{m}_{t}+{\rm clip}_{p,\tau}(\bm{g}_{t}-\bm{m}_{t})\\ \bm{w}_{t+1}=\bm{w}_{t}-\eta_{t}\bm{m}_{t+1}.\end{cases}$
Equation (i) is SGD with momentum where the SPP step size $\tau$ defines the momentum coefficient. On the other hand, (ii) corresponds to clipping, and here the SPP step size determines the clipping threshold. In other words, SGD with momentum is to online mean estimation what clipping is to online median estimation.
Connection to Error Feedback.
Methods such as (15) are also used in the distributed learning setting to improve communication and protect against corrupted nodes (Karimireddy et al., 2021). These methods are often called Error Feedback (EF) methods (Seide et al., 2014; Richtarik et al., 2021) and generally involve an update of the form $\bm{m}_{t+1}=\bm{m}_{t}+\mathcal{C}_{t}(\bm{g}_{t}-\bm{m}_{t})$ where $\mathcal{C}_{t}:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ is a compression operator. Projections onto balls can be seen as a compression, since one can encode the resulting vector in fewer bits (Safaryan et al., 2021). In particular, centered clipping by Karimireddy et al. (2021), and the Clip21 method (Khirirat et al., 2023) are using variations of update (10) in the distributed setting, and are thus ‘secretly estimating’ the geometric median of the gradients returned across all nodes.
SGD with standard clipping.
We will refer to (16) as clipped-SGD. The method has been studied extensively, often for the case $\beta=0$, see Zhang et al. (2020a); Gorbunov et al. (2020); Koloskova et al. (2023); Puchkin et al. (2023) and references therein.
As explained above, when $\beta=0$, then (16) is the same as (10) with cold start (resetting $\bm{m}_{t}=\bm{0}$ in every iteration); but this does not allow for momentum, which is usually used in practice on top of clipping. However, we can derive clipped-SGD (16) also as online SPP for computing a specific robust mean estimator, known as the Catoni-Giulini mean estimator (Catoni & Giulini, 2018; Lugosi & Mendelson, 2019). For this, consider the problem
The solution to (17) is $\mathbb{E}_{\bm{g}\sim\mathcal{G}}\left[\min\{1,\tfrac{c}{\|\bm{g}\|}\}\bm{g}\right]$, which is exactly the Catoni-Giulini mean estimator. Now, one step of SPP (9) applied to (17) gives (16) with $\beta=\tfrac{1}{1+\tau}$.
Outline of remaining paper.
Before comparing these different online gradient estimation methods in experiments, we will discuss convergence guarantees: ideally, we would like to analyze the online method (15). However, when $\mathcal{D}$ is not strongly convex this analysis turns out to be difficult, mainly due to the fact that the distribution of $\bm{g}_{t}$ changes every step. Hence, the next section will analyze the online median estimator in a simplified setting, where we have access to multiple samples each iteration, and can (approximately) compute their sample median. We focus on settings where the noise in $\bm{g}_{t}$ is heavy-tailed, and prove that in such scenarios the sample median can still guarantee convergence (while the sample mean fails).
Finally, our experiments in Section˜6 will show that the online versions of the sample median method (VClip and CClip) show superior performance as well in the heavy-tailed scenario, justifying the online estimation.
5 Theory for the Sample Median Gradient Method
@@
In this section, we showcase that SGD with an (approximate) sample median is robust to heavy-tailed noise, whereas SGD using the sample mean can fail. Consider the following setup: at every iteration $t\geq 0$, we have access to $n$ noisy oracle gradients $\big(\bm{g}_{t}^{(1)},\ldots,\bm{g}_{t}^{(n)}\big)$, where
$\bm{g}_{t}^{(i)}=\nabla\ell(\bm{w}_{t})+\bm{\xi}^{(i)}_{t}(\bm{w}_{t})\quad\mbox{for}\quad i=1,\dots,n.$
Here $\{\bm{\xi}^{(i)}_{t}(\bm{w}_{t})\in\mathbb{R}^{d}\}_{i=1}^{n}$ is a set of i.i.d. noise vectors. Given these sampled gradients, the typical approach of SGD would be to use their sample mean (average) $\bm{g}_{t}=\frac{1}{n}\sum_{i=1}^{n}\bm{g}_{t}^{(i)}$ as update direction. However, Zhang et al. (2020b, Remark 1) show that this fails to converge for any step size under heavy-tailed noise.
Alternatively, to improve robustness, we analyze the update
@@
where $\bm{m}_{t}=\texttt{median}_{\ell_{p},i\in[n]}(\bm{g}_{t}^{(i)})$, and $\bm{e}_{t}\in\mathbb{R}^{d}$ represents a random error we may commit in computing the sample median. We refer to method (18) as Sample Median Gradient Descent (SMGD).
@@
There exists $\sigma_{1},\sigma_{2},\delta_{1},\delta_{2}\geq 0$ such that, for every $\bm{w}\in\mathbb{R}^{d}$, the sample median $\bm{m}=\texttt{median}_{\ell_{p},i\in[n]}(\bm{g}^{(i)})$ verifies
@@
Let $\ell:\mathbb{R}^{d}\to\mathbb{R}$ be $L$-smooth and let Assumption˜5.1 hold with $\delta_{2}\leq\frac{1}{8}$. Take $\eta\leq\frac{1}{8L(1+2\sigma_{2}+2\delta_{2})}$ and consider $\bm{w}_{t}$ a sequence generated by (18) where $\bm{e}_{t}$ verifies Assumption˜5.2. Then, for every $T\geq 1$ it holds
@@
Let $\ell:\mathbb{R}^{d}\to\mathbb{R}$ be convex and $G$-Lipschitz, with $\rm{argmin}$$\ell\neq\emptyset$. Let Assumptions˜5.1 and 5.2 hold. Without loss of generality assume that $\delta_{2}=\sigma_{2}=0$. Then, for $\eta=\frac{1}{(\delta_{1}+\varepsilon)T+\sqrt{T}}$ and every $T\geq 1,$ the iterates $\bm{w}_{t}$ from (18) satisfy
@@
The proofs follow techniques from the analysis of biased SGD (see e.g. Demidovich et al. (2023)) and are provided in Appendix˜A. Let us now discuss our main Assumption˜5.1. Assume for simplicity that $\delta_{2}=\sigma_{2}=0$ (for instance if $\ell$ is Lipschitz), that $\bm{e}_{t}=0$ (the sample median is computed exactly), and consider that $\bm{g}_{t}^{(i)}=\nabla\ell(\bm{w}_{t})+\bm{\xi}_{t}^{(i)}$, where $\bm{\xi}_{t}^{(i)}$ is a noise term. Then, Assumption˜5.1 becomes [^1] See Lemma A.2 in the appendix for a proof.
@@
In other words, we need the second moment of the sample median of $\bm{\xi}_{t}^{(i)}$ to be uniformly bounded. We want to stress that our choice of using the median is crucial when the $\bm{\xi}_{t}^{(i)}$ come from a heavy-tailed distribution, such as a (univariate) standard $\alpha$-stable distribution, with $\alpha\in(1,2)$. Indeed, in this case we know that the variance of the sample median is finite for sufficiently large sample size $n$ (see Bickel (1967, Theorem 2.2) and Nolan (2020, Theorem 1.2)) [^2] We know more: The sample median of a parent distribution with median $\xi$ is asymptotically normal with location $\xi$ (Chu & Hotelling, 1955). For Cauchy distributions ($\alpha=1$), the sample median has a finite variance for $n\geq 5$ (Chu & Hotelling, 1955, Remark 1), while first and second moment of the sample mean are undefined.. On the contrary, taking $\bm{m}_{t}$ as the usual sample mean causes trouble, as the variance of the sample mean is infinite [^3] This is due to the fact that, by definition of $\alpha$-stability, the sample mean of i.i.d sample follows the same $\alpha$-stable distribution (up to affine translation). everywhere. See Fig.˜1 for an illustration of this phenomenon.
This simple discussion illustrates that Assumption˜5.1 can be satisfied under mild conditions on the noise. We exploit this in the next corollary, which shows that the $\ell_{1}$-sample median can even accommodate state-dependent noise. Let us define first our model for such noise.
@@
In the recent literature (see e.g. Zhang et al. (2020b); Puchkin et al. (2023); Sadiev et al. (2023)), a commonly made assumption is that $\mathbb{E}\left[\|\bm{g}-\nabla\ell(\bm{w})\|^{p}\right]\leq\sigma^{2}$ for some $p<2$ where $\sigma^{2}>0$ cannot depend on $\bm{w}$. On the contrary, Assumption˜5.5 can accommodate much stronger noise. To illustrate this, consider $d=1$ and $\ell(\bm{w})=\bm{w}^{2}/2$. Assumption˜5.5 allows the noise model $\bm{\Sigma}(\bm{w})\bm{\zeta}$ with $\bm{\Sigma}(\bm{w})=\sqrt{c_{1}+c_{2}\bm{w}^{2}}$, which results in strong noise, since $\bm{\zeta}$ has infinite variance if $\alpha<2$, and this is amplified by $\bm{\Sigma}(\bm{w})$ when $\bm{w}$ is large. Such strong noise makes vanilla SGD diverge, as we will illustrate in our experiments.
@@
Let $n\geq 3$ be odd and $p=1$. Consider the iterates from (18) with $\bm{e}_{t}=0$, and suppose that $\bm{g}_{t}^{(i)}=\nabla\ell(\bm{w}_{t})+\bm{\Sigma}(\bm{w}_{t})\bm{\zeta}_{t}^{(i)}$ verifies Assumption˜5.5.
@@
    If $\ell$ is $L$-smooth and $\eta=\mathcal{O}(\tfrac{1}{\sqrt{T}})$, then we have that $\min_{0\leq t\leq T-1}\mathbb{E}\left[\|\nabla\ell(\bm{w}_{t})\|^{2}\right]=\mathcal{O}(1/\sqrt{T})$.
@@
    If $\ell$ is convex and Lipschitz and $\eta=1/\sqrt{T}$ then we have that $\mathbb{E}\left[\ell(\bar{\bm{w}}_{T})-\inf\ell\right]=\mathcal{O}(1/\sqrt{T})$, where $\bar{\bm{w}}_{T}$ is defined as in Proposition˜5.4.
@@
This result illustrates how powerful the sample median can be when used in gradient-based optimization. However, SMGD (18) has clear drawbacks: (i) it requires to draw multiple samples per iteration (multiple calls of the oracle), and (ii) we need to compute the sample median up to some tolerance. For instance, the geometric median can not be computed in closed form and there exists a vast body of literature on how to approximate it (Cohen et al., 2016). Moreover, the noise still needs to be symmetric in order for the median to denoise correctly (cf. Assumption˜5.5 and Fig.˜7 (right)).
Nevertheless, our convergence guarantees for SMGD allow us to be optimistic that the strong performance in the heavy-tailed setting will also transfer to the online drop-in replacements (of (18)), which we had previously presented in Section˜4. We investigate this in the following experiments section.
(a) $\alpha=1.1$
(b) $\alpha=1.5$